# QOCO: A Quadratic Objective Conic Optimizer with Custom Solver Generation

### v2

@@ 1 Introduction
Most general-purpose SOCP solvers, such as clarabel [12], cosmo [14], ecos [32], gurobi [13], mosek [33], and scs rely on sparse linear algebra routines. Although sparse linear algebra allows the aforementioned solvers to solve SOCPs regardless of their sparsity structures, it can be quite slow due to the extra overhead of determining where the nonzero elements are before performing floating point operations, which leads to more CPU instructions being issued, more memory accesses, and more cache misses. However, when the sparsity structure of the problem matrices is known beforehand, it is possible to generate a custom solver that implements specialized linear algebra routines tailored to the sparsity structure of the problem. These custom solvers hardcode the exact floating point operations necessary, eliminating the overhead incurred by sparse linear algebra routines, resulting in significantly faster solves.
@@ 1.1 Contribution
This work makes two key contributions: (1) the development of qoco, an open-source C-based solver for quadratic objective SOCPs, and (2) the introduction of qocogen, an open-source custom solver generator for quadratic objective SOCPs which generates a custom solver written in C.
We demonstrate that qoco outperforms most commercial and open-source solvers in terms of both speed and robustness, and qococustom (the name of solvers generated by qocogen) is significantly faster than qoco. Both solvers are easy to use since qoco and qocogen can be called from CVXPY [41, 42] and CVXPYgen [43] respectively, allowing users to formulate optimization problems in a natural way following from math rather than manually converting the problem to the solver-required standard form. Additionally, qoco can be called from C/C++, Matlab, and Python, and qocogen can be called from Python.
Both implement a primal-dual interior point method with Mehrotra’s predictor-corrector [44]. However, they both use a variety of numerical enhancements to quickly and robustly solve the linear system that arises in the step direction computation. Specifically, we use an $LDL^{\top}$ factorization along with the Approximate Minimum Degree (AMD) [45, 46] heuristic to permute the coefficient matrix, minimizing fill-in for the factor $L$. We also apply static and dynamic regularization to the coefficient matrix to ensure that the matrix is invertible and the factorization succeeds. After solving the linear system, we apply iterative refinement to ensure that the original, unregularized system was solved to high accuracy, further enhancing numerical stability and robustness.
qoco and qocogen are available on GitHub at https://github.com/qoco-org.
@@ 1.3 Notation
$(a,b)$ refers to the concatenation of $a\in\mathbb{R}^{a}$ and $b\in\mathbb{R}^{b}$ into a vector in $\mathbb{R}^{a+b}$
The $l$-dimensional non-negative orthant is given by +$\mathbb{R}^{l}_{+}=\{u\in\mathbb{R}^{l}\;|\;u_{i}\geq 0\}$
For a vector $u$ in the $l$-dimensional non-negative orthant, $u_{i}$ refers to the $i^{th}$ element of the vector
The $q$-dimensional second-order cone is given by $\mathcal{Q}^{q}=\{(u_{0},u_{1})\in\mathbb{R}\times\mathbb{R}^{q-1}\;|\;\|u_{1}\|_{2}\leq u_{0}\}$
For a vector $u$ in the $q$-dimensional second-order cone, we partition it as $u=(u_{0},u_{1})$, i.e. $u_{0}\in\mathbb{R}$, $u_{1}\in\mathbb{R}^{q-1}$ and $\|u_{1}\|_{2}\leq u_{0}$.
When $x\in\mathcal{C}_{1}\times\mathcal{C}_{2}\times\cdots\times\mathcal{C}_{K}$, $x_{i}$ refers to the $i^{th}$ subvector which is in cone $\mathcal{C}_{i}$
@@ 2.1 Central path derivation
[…]
@@
If we define $s=h-Gx$ and $z_{i}=-\tau\nabla\phi_{i}(h_{i}-G_{i}x)$ for $i=1,\ldots K$, we can rewrite the KKT conditions as
@@
where the condition $z\in\mathcal{K}$ arises because if $u\in\mathcal{K}$, then $-\nabla\phi(u)\in\mathcal{K}$ [47].
@@ 2.3 Computing search directions
[…]
@@
$\alpha$ $=\sup\{\alpha\in[0,1]\;|\;(s_{k},z_{k})+\alpha(\Delta s_{a},\Delta z_{a})\in\mathcal{K}\times\mathcal{K}\}$
$\rho$ $=\frac{(s_{k}+\alpha\Delta s_{a})^{\top}(z_{k}+\alpha\Delta z_{a})}{s_{k}^{\top}z_{k}}$
$\sigma$ $=\max\{0,\min\{1,\rho\}^{3}\}.$
@@
$r_{x}$ $=Px_{k}+c+A^{\top}y_{k}+G^{\top}z_{k}$
$r_{y}$ $=Ax_{k}-b$
$r_{z}$ $=Gx_{k}+s_{k}-h$
$r_{s}$ $=\lambda_{k}\circ\lambda_{k}+(W_{k}^{-\top}\Delta s_{a})\circ(W_{k}\Delta z_{a})-\sigma\mu_{k}\boldsymbol{e}$
This direction moves the next iterate closer to $(x^{*}(\sigma\mu_{k}),s^{*}(\sigma\mu_{k}),y^{*}(\sigma\mu_{k}),z^{*}(\sigma\mu_{k}))$, a point on the central path where the duality measure is reduced by a factor of $\sigma$. Note that the residual vector $r_{s}$ has two additional terms when compared to Equation (20d): $(W_{k}^{-\top}\Delta s_{a})\circ(W_{k}\Delta z_{a})$ and $-\sigma\mu_{k}\boldsymbol{e}$. The former is to correct for linearization error in Newton’s method, and the latter is to maintain proximity to the central path and arises when applying Newton’s method to Equation (9d) with $\tau=\sigma\mu_{k}$.
@@ 2.4 Initialization
We then set $s=-z$. This results in the tuple $(x,s,y,z)$ satisfying Equations (9a) - (9c), but $s$ and $z$ may not necessarily satisfy Equation (9e).
@@ 4.2 Code Generation
qocogen is written is Python and example code for generating qococustom for a problem family defined by Equation (31) is given by Listing 2. The generated solver qococustom is customized to the sparsity pattern of the matrices $P,A,G$ which are passed to the function qocogen.generate_solver() as sparse scipy matrices.
[…]
@@ 5 Numerical Results
We benchmarked qoco and qococustom against several conic interior-point solvers: the open-source solvers ecos [32] and clarabel [12], the commercial solvers gurobi [13] and mosek [33], and against the academically licensed custom solver generator cvxgen which generates a primal-dual interior point method for quadratic programs. We did not test against the custom solver generator bsocp since it is not academically licensed. For all solvers, we used their default settings but set the tolerances $\epsilon_{abs}=\epsilon_{rel}=10^{-7}$, and disabled the presolve features of gurobi and mosek to ensure all solvers are given the same problem formulation.
We evaluated our solvers on benchmark problems spanning multiple fields and a range of problem sizes. In our experiments, we define problem size as the total number of nonzero elements in $A$, $G$, and in the upper half of $P$. For all benchmark problems, we tested qoco, but only tested qococustom for the smaller problems, because code generation and compile times become prohibitively long for larger problem instances. We also evaluated qoco on the difficult Maros–Mészáros problems, a repository of 138 challenging quadratic programs (QPs) [61]. To evaluate the solvers’ performance, we used performance profiles [62] and the shifted geometric mean, both of which are frequently used to assess solver performance [12, 32, 56, 14, 63].
All numerical results were generated on a desktop computer with an AMD Ryzen 9 7950X3D processor running at 4.2 GHz and with 128 GB of RAM. Our benchmarks are written in Python and we used CVXPY [41, 42] to interface with the solvers. Our benchmarks are publicly available [64].
@@
where $t_{s,p}$ is the runtime of solver $s$ on problem $p$ and $k=1$ is the shift. If the solver does not find an optimal solution, then the runtime is set to the maximum allowable runtime (10 seconds for benchmark problems and 1200 seconds for the Maros–Mészáros problems).
@@ 5.1 Benchmark Problems
We solved optimization problems from five problem classes: robust Kalman filtering, group lasso regression [65], the losslessly convexified powered-descent guidance problem [37], Markowitz portfolio optimization [3], and the oscillating masses control problem [66]. The first three of these are SOCPs and the last two are QPs. For each of the five classes, we considered 10 different problem sizes, generating 20 unique problem instances per size, for a total of 1000 distinct optimization problems. Details on the problems we solved can be found in Appendix C. For these problems, we plotted runtime as a function of problem size in addition to reporting performance profiles and shifted geometric means.
For each solver, we solved each optimization problem 100 times and recorded the minimum runtime, which includes both setup and solve time. Since the computer runs various background processes that can artificially inflate execution time, we conducted 100 runs and reported the minimum execution time, thereby providing a more accurate measure of the solver’s performance.
Although the oscillating masses and portfolio optimization problems are QPs and could be solved with cvxgen, we limited testing to the two smaller oscillating masses problems. For larger instances, cvxgen was unable to generate code due to the size of the problems. For the portfolio optimization problems, we opted not to test cvxgen due to the sparsity pattern of the factor matrix $F$ in Problem 34. To specify the sparsity pattern of $F$ in cvxgen, we must manually provide the locations of nonzero elements into the web interface, which would have been prohibitively expensive, since $F$ has hundreds of nonzero elements. In contrast, to generate code with qocogen, we pass in problem data as sparse scipy matrices rather than manually specifying sparsity patterns. We could have specified $F$ as a dense matrix and only populated the nonzero elements for cvxgen, this would severely hinder its performance and result in an unfair comparison.
@@
Figures 4, 5, and 6 show that qoco is the fastest generic solver on our benchmark problems. Additionally, qococustom is the fastest solver by a significant margin and is as performant as cvxgen on the $40$ smallest oscillating mass problems. This latter result is unsurprising, as qococustom uses an identical algorithm to cvxgen when solving QPs and both utilize custom linear algebra routines.
@@ 5.2 Maros–Mészáros problems
The Maros–Mészáros problems are a set of 138 QPs that include a wide range of problem sizes (from $3$ to around $300,000$ variables) and contain very difficult problems that have ill-conditioning and poor scaling [61]. These properties make the Maros–Mészáros problems a good test set to assess the robustness of solvers. We test qoco but not qococustom on these problems since many problems are far too large to generate code for. Additionally, we ran each problem once rather than the 100 runs we did for the benchmark problems since many of the Maros–Mészáros problems are extremely large, and running them 100 times per solver would be extremely cumbersome.
Figure 7 shows that the proprietary solver Gurobi successfully solved the largest fraction of the problems followed by qoco, clarabel, mosek, and ecos. We can also see that for the majority of the problems, qoco was the fastest solver, whereas mosek and ecos tended to be the slowest. The latter result is unsurprising as mosek and ecos can only handle linear objectives and the reformulation of the quadratic objective into a second-order cone likely introduced significant fill-in, slowing these solvers down. We also observed that for many of the largest problems, Gurobi was the fastest solver, due to multithreading in its matrix factorization.
@@
Shifted GM 2.0 3.0 27.4 1.0 3.8
Failure Rate (%) 6.5 8.7 31.9 2.9 9.4

### v3

@@ 1 Introduction
$\mathcal{K}:=\mathcal{C}_{1}\times\mathcal{C}_{2}\times\cdots\times\mathcal{C}_{K},$
@@
Most general-purpose SOCP solvers, such as clarabel [12], cosmo [14], ecos [32], gurobi [13], mosek [33], and scs rely on sparse linear algebra routines. Although sparse linear algebra allows the aforementioned solvers to solve SOCPs regardless of their sparsity structures, it can be quite slow due to the extra overhead of determining where the nonzero elements are before performing floating point operations, which leads to more CPU instructions being issued, more memory accesses, and more cache misses. However, when the sparsity structure of the problem matrices is known beforehand, it is possible to generate a custom solver that implements customized linear algebra routines tailored to the sparsity structure of the problem. Specifically, in sparse linear algebra matrices are stored as three arrays: a row/column index array, column/row pointer array, and data array. Before performing a floating point operation on a nonzero element in the matrix, sparse linear algebra routines must first index into the pointer and index arrays to determine the location of the element in the data array. In customized linear algebra, the index and pointer arrays do not exist, and we directly index into the data array, eliminating the overhead of sparse linear algebra, resulting in significantly faster operations. We discuss this further in Section 4.1.
@@
A seemingly related but distinct tool is cvxpygen [41]. It leverages cvxpy to parse the optimization problem to the standard form for solvers and generates C code for updating problem data and retrieving solutions. cvxpygen then wraps a backend solver without (or with very little) modification to its source code. Because it customizes only the parsing and solution retrieval while relying entirely on an existing solver, we do not classify cvxpygen as a custom solver generator.
@@ 1.1 Contribution
This work makes two key contributions: (1) the development of qoco, an open-source C-based solver for quadratic objective SOCPs, and (2) the introduction of qocogen, an open-source custom solver generator for quadratic objective SOCPs which generates a custom solver written in C  [^3] Both are available on GitHub at https://github.com/qoco-org.
We demonstrate that qoco outperforms most commercial and open-source solvers in terms of both speed and robustness, and qococustom (the name of solvers generated by qocogen) is significantly faster than qoco. Both solvers are easy to use since qoco and qocogen can be called from cvxpy [42, 43] and cvxpygen [41] respectively, allowing users to formulate optimization problems in a natural way following from math rather than manually converting the problem to the solver-required standard form. Additionally, qoco can be called from C/C++, Matlab, and Python, and qocogen can be called from Python. Both implement a primal-dual interior point method with Mehrotra’s predictor-corrector [44]. However, they both use a variety of numerical enhancements to quickly and robustly solve the linear system that arises in the step direction computation. Specifically, we use an $LDL^{\top}$ factorization along with the Approximate Minimum Degree (AMD) [45, 46] heuristic to permute the coefficient matrix, minimizing fill-in for the factor $L$. We also apply static and dynamic regularization to the coefficient matrix to ensure that the matrix is invertible and the factorization succeeds. After solving the linear system, we apply iterative refinement to ensure that the original, unregularized system was solved to high accuracy, further enhancing numerical stability and robustness.
@@ 1.3 Notation
We use $(a,b)$ to denote the concatenation of vectors $a\in\mathbb{R}^{a}$ and $b\in\mathbb{R}^{b}$, resulting in a vector in $\mathbb{R}^{a+b}$. The non-negative orthant in $\mathbb{R}^{l}$ is denoted by $\mathbb{R}^{l}_{+}=\{u\in\mathbb{R}^{l}\mid u_{i}\geq 0\}$, and for any $u\in\mathbb{R}^{l}_{+}$, we use $u_{i}$ to refer to its $i^{\text{th}}$ element. The $q$-dimensional second-order cone is defined as $\mathcal{Q}^{q}=\{(u_{0},u_{1})\in\mathbb{R}\times\mathbb{R}^{q-1}\mid\|u_{1}\|_{2}\leq u_{0}\}$, where any $u\in\mathcal{Q}^{q}$ is written as $u=(u_{0},u_{1})$ with $u_{0}\in\mathbb{R}$ and $u_{1}\in\mathbb{R}^{q-1}$. Finally, when a vector $x$ lies in the Cartesian product of cones, i.e., $x\in\mathcal{C}_{1}\times\mathcal{C}_{2}\times\cdots\times\mathcal{C}_{K}$, we denote $x_{i}$ as its $i^{\text{th}}$ subvector belonging to cone $\mathcal{C}_{i}$.
@@ 2.1 Central path derivation
$\underset{x}{\text{minimize}}$ $\frac{1}{2}x^{\top}Px+c^{\top}x+\mathcal{I}_{\mathcal{K}}(h-Gx)$
@@
If we define s=hGxs=h-Gx and $z_{i}=-\tau\nabla\phi_{i}(h_{i}-G_{i}x)$ for $i=1,\ldots K$, we can rewrite the KKT conditions as
@@
where the condition $z\in\mathcal{K}$ arises because if $u\in\mathcal{K}$, then $\nabla\phi(u)\in\mathcal{K}$ [47].
@@ 2.3 Computing search directions
where the operator $\backslash$ represents the inverse of the Jordan product $\circ$, meaning that $x\backslash(x\circ y)=y$.
@@
$\begin{split}\alpha&=\sup\{\alpha\in[0,1]\;|\;(s_{k},z_{k})+\alpha(\Delta s_{a},\Delta z_{a})\in\mathcal{K}\times\mathcal{K}\}\\ \rho&=\frac{(s_{k}+\alpha\Delta s_{a})^{\top}(z_{k}+\alpha\Delta z_{a})}{s_{k}^{\top}z_{k}}\\ \sigma&=\max\{0,\min\{1,\rho\}^{3}\}.\end{split}$
@@
$r_{x}$ $=Px_{k}+c+A^{\top}y_{k}+G^{\top}z_{k}$ (21a)
$r_{y}$ $=Ax_{k}-b$ (21b)
$r_{z}$ $=Gx_{k}+s_{k}-h$ (21c)
$r_{s}$ $=\lambda_{k}\circ\lambda_{k}+(W_{k}^{-\top}\Delta s_{a})\circ(W_{k}\Delta z_{a})-\sigma\mu_{k}\boldsymbol{e}$ (21d)
This direction moves the next iterate closer to $(x^{*}(\sigma\mu_{k}),s^{*}(\sigma\mu_{k}),y^{*}(\sigma\mu_{k}),z^{*}(\sigma\mu_{k}))$, a point on the central path where the duality measure is reduced by a factor of $\sigma$. Note that the residual vector $r_{s}$ has two additional terms when compared to Equation (20d): $(W_{k}^{-\top}\Delta s_{a})\circ(W_{k}\Delta z_{a})$ and $\sigma\mu_{k}\boldsymbol{e}$. The former is to correct for linearization error in Newton’s method, and the latter is to maintain proximity to the central path and arises when applying Newton’s method to Equation (9d) with $\tau=\sigma\mu_{k}$.
@@ 2.4 Initialization
We then set s=zs=-z. This results in the tuple $(x,s,y,z)$ satisfying Equations (9a) - (9c), but $s$ and $z$ may not necessarily satisfy Equation (9e).
@@ 4.2 Code generation
qocogen is written in Python and example code for generating qococustom for a problem family defined by Equation (30) is given by Listing 2. The generated solver qococustom is customized to the sparsity pattern of the matrices $P,A,G$ which are passed to the function qocogen.generate_solver() as sparse scipy matrices.
$\begin{array}[]{ll}\underset{x}{\text{minimize}}&x_{1}^{2}+x_{2}^{2}+x_{3}^{2}+x_{4}\\ \mbox{subject to}&x_{1}+x_{2}=1\\ &x_{2}+x_{3}=1\\ &x_{1}\geq 0\\ &\sqrt{x_{3}^{2}+x_{4}^{2}}\leq x_{2}\end{array}$ (30)
@@ 5 Numerical results
We benchmark qoco and qococustom against several conic interior-point solvers  [^4] Our benchmarks are publicly available at https://github.com/qoco-org/qoco-benchmarks: the open-source solvers ecos [32] and clarabel [12], the commercial solvers gurobi [13] and mosek [33], and against the academically licensed custom solver generator cvxgen which generates a primal-dual interior point method for quadratic programs. We do not test against the custom solver generator bsocp since it is not academically licensed. For all solvers, we use their default settings but set the tolerances $\epsilon_{abs}=\epsilon_{rel}=10^{-7}$.
In our experiments, we define problem size as the total number of nonzero elements in $A$, $G$, and in the upper half of $P$. We evaluate our solvers on a set of benchmark problems we developed, a set of model-predictive control problems from [61], the Maros–Mészáros problems [62], and least-squares problems derived from the SuiteSparse repository [63]. To evaluate the solvers’ performance, we use performance profiles [64] and the shifted geometric mean, both of which are frequently used to assess solver performance [12, 32, 56, 14, 65].
All numerical results were generated on a desktop computer with an AMD Ryzen 9 7950X3D processor running at 4.2 GHz and with 128 GB of RAM. Our benchmarks are written in Python and we use cvxpy [42, 43] to interface with the solvers.
@@
where $t_{s,p}$ is the runtime of solver $s$ on problem $p$ and $k=1$ is the shift. If the solver does not find an optimal solution, then the runtime is set to 10 seconds for benchmark problems, 0.5 seconds for the MPC problems, 1200 seconds for the Maros–Mészáros problems, and 600 seconds for the SuiteSparse least-squares problems. These times were chosen to be roughly an order of magnitude higher than the slowest successful solve on the respective problem set.
@@ 5.1 Benchmark problems
We solve optimization problems from five problem classes: robust Kalman filtering, group lasso regression [66], the losslessly convexified powered-descent guidance problem [37], Markowitz portfolio optimization [3], and the oscillating masses control problem [67]. The first three of these are SOCPs and the last two are QPs. For each of the five classes, we consider 10 different problem sizes, generating 20 unique problem instances per size, for a total of 1000 distinct optimization problems. Details on the problems we solve can be found in Appendix C. For these problems, we plot runtime as a function of problem size in addition to reporting performance profiles and shifted geometric means. We test qoco on all problems, but only test qococustom on the smaller problems, as code generation and compile times become prohibitively long for the larger instances.
For each solver, we solve each optimization problem 100 times and record the minimum runtime, which includes both setup and solve time. Since the computer runs various background processes that can artificially inflate execution time, we conduct 100 runs and report the minimum execution time, thereby providing a more accurate measure of the solver’s performance.
Although the oscillating masses and portfolio optimization problems are QPs and can be solved with cvxgen, we limit testing to the two smaller oscillating masses problems. For larger instances, cvxgen was unable to generate code due to the size of the problems. For the portfolio optimization problems, we do not test cvxgen due to the sparsity pattern of the factor matrix $F$ in Problem 33. To specify the sparsity pattern of $F$ in cvxgen, we must manually provide the locations of nonzero elements into the web interface, which would have been prohibitively expensive, since $F$ has hundreds of nonzero elements. In contrast, to generate code with qocogen, we pass in problem data as sparse scipy matrices rather than manually specifying sparsity patterns. Although we can specify $F$ as a dense matrix and only populate the nonzero elements for cvxgen, this would severely hinder its performance and result in an unfair comparison.
@@
Figures 4, 5, and 6 show that qococustom is the fastest solver on our benchmark problems by a significant margin and is as performant as cvxgen on the $40$ smallest oscillating mass problems. This latter result is unsurprising, as qococustom uses an identical algorithm to cvxgen when solving QPs and both utilize custom linear algebra routines. Additionally, qoco is the fastest open-source generic solver on all of the problems, although it is slower than Mosek on the group lasso problems and slower than Gurobi on some of the largest portfolio optimization problems. However, we notice that if we disable presolve for Mosek and Gurobi, qoco is faster on these problems as well.
@@
5.2 Model-predictive control problems
We consider a set of 64 model-predictive control problems taken from [61] and also used by [12]. Note that we exclude the eight “nonlinear chain” problems as they were too large to generate custom solvers for. Similarly to the benchmark problems, we solve each MPC problem 100 times and record the minimum runtime. These problems take the following form
[…]
Figure 7 shows that clarabel, gurobi, and mosek solve all problems successfully, qoco and qococustom each fail on only one, and ecos fails on seven. Among the 63 problems solved by both qoco and qococustom, qococustom is the fastest solver by a wide margin, followed by qoco.
(a) Relative performance profile
(b) Absolute performance profile
QOCO Custom QOCO Clarabel ECOS Gurobi Mosek
Shifted GM 4.4 4.5 1.0 31.4 1.7 2.5
Failure Rate (%) 1.6 1.6 0.0 10.9 0.0 0.0
(c) Shifted geometric means and failure rates
Figure 7: Performance profiles for model-predictive control problems
@@ 5.3 Maros–Mészáros problems
The Maros–Mészáros problems are a set of 138 challenging QPs that include a wide range of problem sizes (from $3$ to around $300,000$ variables) and contain very difficult problems that have ill-conditioning and poor scaling [62]. These properties make the Maros–Mészáros problems a good test set to assess the robustness of solvers. We test qoco but not qococustom on these problems since many problems are far too large to generate code for. Additionally, we run each problem once rather than the 100 runs we do for the benchmark and MPC problems since many of the Maros–Mészáros problems are extremely large, and running them 100 times per solver is extremely cumbersome.
Figure 8 shows that the proprietary solver Gurobi successfully solved the largest fraction of the problems followed by qoco, clarabel, mosek, and ecos. We can also see that for the majority of the problems, qoco was the fastest solver, whereas mosek and ecos tended to be the slowest. The latter result is unsurprising as mosek and ecos can only handle linear objectives and the reformulation of the quadratic objective into a second-order cone likely introduces significant fill-in, slowing these solvers down. We also observe that for many of the largest problems, Gurobi is the fastest solver, due to multithreading in its matrix factorization.
@@
Shifted GM 2.4 3.6 33.9 1.0 4.1
Failure Rate (%) 6.5 8.7 31.9 2.9 8.7
@@
5.4 SuiteSparse least-squares problems
Finally, we consider a set of 23 least-squares problems $Ax\approx b$, where the matrix $A\in\mathbb{R}^{m\times n}$ is drawn from the SuiteSparse Matrix Collection [63]. Following [56] and [12], we get an approximate solution to each least-squares problem by solving a Huber regression and lasso regression problem, which can both be formulated as constrained quadratic programs [51]. These optimization problems are quite large with some having over a million optimization variables, so we do not test qococustom due to the prohibitively long code-generation time. Additionally, we run each problem once rather than 100 times.
The Huber regression problem is given by
[…]
where $a_{i}^{\top}$ is the $i^{th}$ row of $A$ and $\phi:\mathbb{R}\to\mathbb{R}$ is defined as
$\phi(z)=\begin{cases}z^{2}&\text{if }|z|\leq\delta\\ \delta(2|z|-\delta)&\text{if }|z|>\delta\end{cases},$
where we use $\delta=1$.
The lasso regression problem is given by
[…]
where we set $\lambda=\|A^{\top}b\|_{\infty}$.
Figure 9 shows that qoco and clarabel successfully solve all problems, with qoco being the fastest on most instances. As with the Maros–Mészáros problems, we observe that Gurobi is the fastest on some of the largest problems, likely due to it multithreaded matrix factorization.
(a) Relative performance profile
(b) Absolute performance profile
QOCO Clarabel ECOS Gurobi Mosek
Shifted GM 1.0 1.2 7.6 1.6 1.2
Failure Rate (%) 0.0 0.0 23.9 8.7 2.2
(c) Shifted geometric means and failure rates
Figure 9: Performance profiles for SuiteSparse least-squares problems